A string of length 'L' and mass 'M' is lyin on a horizontal table. a force'F' is applied at one of its ends. tension in the string at a distance x from which force is applied is.....................

Grade:11

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Nagarjun,

F = Ma

and T = (xM/L)a  at distance x

so, T = xF/L

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

Lav Kumar Yadav
13 years ago
mass of string =M length of string=L. T=ma where m is mass of L-x distance of string from other end T=M/l *L-x *a put a=F/M and get your ans